# Inverse semigroups of separated graphs and associated algebras

### v2

@@ Abstract.
The first author was partially supported by the Spanish State Research Agency (grants No. PID2020-113047GB-I00/AEI/10.13039/501100011033 and CEX2020-001084-M), and by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No. 2017-SGR-1725). The second author was supported by CNPq/CAPES/Humboldt/Fapesc, and the third author by CNPq - 402924/2022-3.
@@ 1. Introduction
The purpose of this paper is to introduce the inverse semigroup $\mathcal{S}(E,C)$ of a separated graph $(E,C)$ and to develop Exel’s program [Exel:Inverse_combinatorial] in order to describe the structure of the tight algebras associated with this inverse semigroup. To achieve this goal, we will examine the internal structure of the semigroup $\mathcal{S}(E,C)$, obtaining a detailed description of its semilattice of idempotents, which is a crucial component in our analysis. The inverse semigroup $\mathcal{S}(E,C)$ generalizes the graph inverse semigroup $\mathcal{S}(E)$ first introduced by Ash and Hall in [AshHall], which has been widely analized by several authors, see e.g. [JonesLawson, LaLonde-ConditionK, meakin-wang-2021, meakin-milan-wang-2021, mesyan-mitchell-2016].
@@ 2.4. Strongly E*-unitary inverse semigroups and partial actions
If $\Gamma$ is a group acting partially on a semilattice $\mathcal{E}$ via partial isomorphisms $\theta_{g}\colon D_{g^{-1}}\to D_{g}$ between ideals $D_{g}\mathrel{\trianglelefteq}\mathcal{E}$, then we define its semidirect product as $\mathcal{E}\rtimes_{\theta}\Gamma=\{(e,g):e\in D_{g},g\in\Gamma\}$. We write elements of $\mathcal{E}\rtimes_{\theta}\Gamma$ as $e\delta_{g}$ with $e\in D_{g}$, $g\in G$. This becomes an inverse semigroup when endowed with the following product and inverse:
@@
Let $S$ be a strongly $E^{*}$-inverse semigroup with 00, and let $\pi\colon S^{\times}\to\Gamma$ be an idempotent pure partial homomorphism. Then $\pi$ induces a partial action $\theta=(\theta_{g}\colon D_{g^{-1}}\to D_{g})_{g\in\Gamma}$ of $\Gamma$ on $\mathcal{E}=\mathcal{E}(S)$ and a natural isomorphism $S\cong\mathcal{E}\rtimes_{\theta}^{r}\Gamma$. The ideal $D_{g}$ is defined as the set of all the elements $ss^{-1}$, where $s\in S^{\times}$ and $\pi(s)=g$. The partial action is defined by $\theta_{g}(s^{-1}s)=ss^{-1}$ for $s^{-1}s\in D_{g^{-1}}$, and the isomorphism $\mathcal{E}\rtimes^{r}_{\theta}\Gamma\to S$ sends $e\delta_{g}$ to $s$, where $e=ss^{-1}$ and $\theta(s)=g$.
@@
More precisely, let $S$ is a strongly $E^{*}$-inverse semigroup, with an idempotent pure partial homomorphism $\pi\colon S^{\times}\to\Gamma$. Observe that $S$ is the ideal quotient of the $E$-unitary inverse semigroup
@@ 4. Munn trees
Note that $\Gamma_{E}$ is a forest, that is, a disjoint union of trees. Each connected component $C$ of $\Gamma_{E}$ contains a unique vertex $v$, and $C$ consists of all the reduced paths $p$ such that $s(p)=v$. We will denote by $C_{v}$ the connected component of $v$. Observe that for each $v\in E^{0}$, the connected component $C_{v}$ is isomorphic to a certain subtree of the Cayley graph $\Gamma$ of the free group $\mathbb{F}$ on $E^{1}$, namely the subtree consisting in replacing the vertex $v$ by $1\in\mathbb{F}$, and identiying each other vertex of $C_{v}$ with the corresponding vertex in $\Gamma$.
There is natural partial action by left translations of $\mathfrak{F_{gr}}(E)$ on the graph $\Gamma_{E}$.
Let $(E,C)$ be a separated graph. Observe that, for $v\in E^{0}$, elements in $\mathcal{X}(v)$ can be interpreted as certain finite subtrees of $C_{v}$ containing the root $v$ of the connected component $C_{v}$. Concretely, the finite subtrees $T$ which belong to $\mathcal{X}(v)$ are characterized by the property that the geodesic path between any two vertices of $T$ is a $C$-separated path (see Remark 3.8).
@@
where $T_{1}^{\prime},T_{2}^{\prime}\in\mathcal{E}$ are representatives of $T_{1}$ and $T_{2}$ respectively, such that $g_{1}\in T_{1}^{\prime}$ and $g_{2}\in T_{2}^{\prime}$. The inverse is given by $(T,g)^{-1}=(g^{-1}\cdot T^{\prime},g^{-1})$, where $T^{\prime}$ is a representative of $T$ such that $g\in T^{\prime}$.
@@
Let $(E,C)$ be a separated graph and let $\mathcal{E}$ be the semilattice of idempotents of $\mathcal{S}(E,C)$. Then there exists an isomorphism $\mathcal{Y}\cong\mathcal{E}$ given by
@@
where $I_{0}$ is the unique maximal element of $[I]$, which is characterized by the property that each maximal element of $I_{0}$ does not end in $E^{-1}$.
@@
We now analize Examples 3.3 and 3.4 in the light of the Munn tree picture.
@@ 5. The tight spectrum
Before we continue with our analysis of (inverse) semigroups of separated graphs, let us discard a trivial part of them, namely the isolated vertices, that is, the vertices $v\in E^{0}$ with $s^{-1}(v)=r^{-1}(v)=\emptyset$. If $E^{0}_{\mathrm{iso}}$ denotes the set of isolated vertices, notice that these are idempotent elements of both semigroups $S=\tilde{\mathcal{S}}(E,C)$ or $S=\mathcal{S}(E,C)$ that are orthogonal to every other elements. This means that $S$ contains a (commutative) inverse subsemigroup $E^{0}_{\mathrm{iso}}\cup\{0\}$ consisting of pairwise orthogonal idempotents that do not interact with the other non-zero elements of the semigroup. This “isolated part” of $S$ will not be interesting in any respect, as for instance in the algebras we are going to consider later, this only contributes to a trivial commutative direct sum of some copies of the underlying ground field (or ring). Hence there is no loss of generality in assuming $E^{0}_{\mathrm{iso}}=\emptyset$, so we postulate, from now on, the following:
@@
With the above notation we have
@@
    The set $\mathcal{F}$ of filters of $\mathcal{Y}$ is order-isomorphic with $\mathfrak{F}$.
@@
    The set of ultrafilters of $\mathcal{Y}$ is isomorphic to $\mathfrak{U}$.
@@
We finally show that $Z$ is maximal in $\mathfrak{F}$, equivalenty in $\mathfrak{F}(v)$. By way of contradiction suppose that $Z\subsetneqq\tilde{Z}$ for some $\tilde{Z}\in\mathfrak{F}(v)$. Take $g\in\tilde{Z}\setminus Z$, and let $h$ be the largest prefix of $g$ such that $h\in Z$, so that $hy\in\tilde{Z}\setminus Z$ for some $y\in E^{1}\cup E^{-1}$. Suppose first that $y\in E^{1}$. Then $y\in X$ for a unique $X\in C_{r(h)}=C_{s(y)}$. On the other hand, $h\in Z^{\prime}$ and thus $Z^{\prime}_{h}$ is a maximal local configuration. It follows that there is a unique $x\in X\cap Z^{\prime}_{h}$, so that $hx\in Z^{\prime}$. Hence $hx\in Z$, because it belongs to $Z^{\prime}$ and ends in $E^{1}$. (Observe that $h$ cannot end in $x^{-1}$, because $hy$ is a reduced word that belongs to $\tilde{Z}$, and so it belongs to $\mathbb{F}_{C}(v)$; hence $hx$ is reduced.) If $x=y$ then $hy\in Z$ which is not the case, hence $x\neq y$. But then, since $Z\subseteq\tilde{Z}$, we have $hx\in\tilde{Z}$. Since also $hy\in\tilde{Z}$, we obtain a contradiction, because $\tilde{Z}$ is $C$-compatible, and $hx$ and $hy$ are $C$-incompatible. So in any case we get a contradiction, and thus $y\in E^{-1}$. Write $y=e_{1}^{-1}$, where $e_{1}\in E^{1}$. We have $he_{1}^{-1}\in\tilde{Z}$. Moreover since $Z^{\prime}_{h}$ is a maximal local configuration, $e_{1}^{-1}\in Z^{\prime}_{h}$, which implies that $he_{1}^{-1}\in Z^{\prime}$. Now since $\tilde{Z}\in\mathfrak{F}(v)$, we can extend $he_{1}^{-1}$ to a reduced word ending in $E^{1}$, so that there are $e_{2},\dots,e_{n},d\in E^{1}$ such that
@@
    $\mathcal{N}(v)=\{x_{1}^{-1}x_{2}^{-1}\cdots x_{n}^{-1}y\in\mathbb{F}_{C}(v):n\geq 0,x_{1},\dots,x_{n},y\in E^{1}\}.$
@@
    $\mathcal{Z}(I\setminus F)=\mathcal{U}(\{I\},\{I\cup\{gg^{\prime}\}^{\downarrow}:gg^{\prime}\in F\}),$
@@
Let $(E,C)$ be a separated graph. Then the space $\mathfrak{U}$ of ultrafilters on the semilattice of idempotents of $\mathcal{S}(E,C)$ is closed in the space of all filters if and only if $(E,C)$ is finitely separated.
@@ 6.1. Leavitt path algebras and tight groupoids
of $Z$ are mutually orthogonal. Since all then sets $Z(f)$ are non-empty, $Z$ is finite, and $Y_{0}$ is infinite, we obtain a contradiction. We have thus shown that any $z\in Z$ such that $I_{z}$ contains an element $\alpha\in I_{z}\setminus I$ such that $\alpha_{I_{z}\setminus I}=\gamma_{0}e\gamma_{1}$, where $e\in Y$ and $|Y|=\infty$ is redundant. Therefore we can assume that all elements of $Z$ have the property that, for each $\alpha=\alpha_{I}\alpha_{I_{z}\setminus I}\in I_{z}$, all edges appearing (with positive exponent) in the path $\alpha_{I_{z}\setminus I}$ belong to some $X\in C^{\mathrm{fin}}$.
@@ Acknowledgments
The authors are very grateful to Ben Steinberg for his many comments and suggestions on an earlier version of the paper. In particular, the approach to Theorem 4.5 using Munn trees and the free inverse semigroup $\mathfrak{F_{is}}(E)$ is due to him.

### v3

@@ Abstract.
The first author was partially supported by the Spanish State Research Agency (grants No. PID2020-113047GB-I00/AEI/10.13039/501100011033, PID2023-147110NB-I00, and CEX2020-001084-M), and by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No. 2021-SGR-01015). The second author was supported by CNPq, CAPES/Humboldt and FAPESC, and the third author by CNPq - 402924/2022-3.
@@ 1. Introduction
The purpose of this paper is to introduce the inverse semigroup $\mathcal{S}(E,C)$ of a separated graph $(E,C)$ and to develop Exel’s program [Exel:Inverse_combinatorial] in order to describe the structure of the tight algebras associated with this inverse semigroup. To achieve this goal, we will examine the internal structure of the semigroup $\mathcal{S}(E,C)$, obtaining a detailed description of its semilattice of idempotents, which is a crucial component in our analysis. The inverse semigroup $\mathcal{S}(E,C)$ generalizes the graph inverse semigroup $\mathcal{S}(E)$ first introduced by Ash and Hall in [AshHall], which has been widely analyzed by several authors, see e.g. [JonesLawson, LaLonde-ConditionK, meakin-wang-2021, meakin-milan-wang-2021, mesyan-mitchell-2016].
@@
The structural results developed in this work not only clarify the internal properties of $\mathcal{S}(E,C)$, but also provide a natural framework to analyze the associated algebras – such as Leavitt path algebras and graph C-algebras – via partial crossed product decompositions. This perspective connects the combinatorial nature of separated graphs with the dynamics of étale groupoids, offering tools to investigate key properties of the associated algebras, including simplicity, pure infiniteness, and K-theory. Such decomposition techniques have already proven effective in the study of these algebras in earlier works (e.g., [Ara-Exel:Dynamical_systems, AraLolk, Lolk:tame, Lolk:nuclearity]), and continue to play a central role in ongoing developments. While a full exploration of these implications lies beyond the scope of the present paper, we expect that the framework presented here will support and inspire further investigations in this direction.
@@ 2.2. Separated graphs
If all the sets in $C$ are finite, we say that $(E,C)$ is a finitely separated graph.
@@ 2.4. Strongly E*-unitary inverse semigroups and partial actions
If $\Gamma$ is a group acting partially on a semilattice $\mathcal{E}$ via partial isomorphisms $\theta_{g}\colon D_{g^{-1}}\to D_{g}$ between ideals $D_{g}\mathrel{\trianglelefteq}\mathcal{E}$, then we define its semidirect product as $\mathcal{E}\rtimes_{\theta}\Gamma=\{(e,g):e\in D_{g},g\in\Gamma\}$. We write elements of $\mathcal{E}\rtimes_{\theta}\Gamma$ as $e\delta_{g}$ with $e\in D_{g}$, $g\in\Gamma$. This becomes an inverse semigroup when endowed with the following product and inverse:
@@
Let $S$ be a strongly $E^{*}$-inverse semigroup with 00, and let $\pi\colon S^{\times}\to\Gamma$ be an idempotent pure partial homomorphism. Then $\pi$ induces a partial action $\theta=(\theta_{g}\colon D_{g^{-1}}\to D_{g})_{g\in\Gamma}$ of $\Gamma$ on $\mathcal{E}=\mathcal{E}(S)$ and a natural isomorphism $S\cong\mathcal{E}\rtimes_{\theta}^{r}\Gamma$. The ideal $D_{g}$ is defined as the set of all the elements $ss^{-1}$, where $s\in S^{\times}$ and $\pi(s)=g$, together with 00. The partial action is defined by $\theta_{g}(s^{-1}s)=ss^{-1}$ for $s^{-1}s\in D_{g^{-1}}$, and the isomorphism $\mathcal{E}\rtimes^{r}_{\theta}\Gamma\to S$ sends $e\delta_{g}$ to $s$, where $e=ss^{-1}$ and $\pi(s)=g$.
@@
More precisely, let $S$ be a strongly $E^{*}$-inverse semigroup, with an idempotent pure partial homomorphism $\pi\colon S^{\times}\to\Gamma$. Observe that $S$ is the ideal quotient of the $E$-unitary inverse semigroup
@@ 4. Munn trees
Note that $\Gamma_{E}$ is a forest, that is, a disjoint union of trees. Each connected component $K$ of $\Gamma_{E}$ contains a unique vertex $v$, and $K$ consists of all the reduced paths $p$ such that $s(p)=v$. We will denote by $K_{v}$ the connected component of $v$. Observe that for each $v\in E^{0}$, the connected component $K_{v}$ is isomorphic to a certain subtree of the Cayley graph $\Gamma$ of the free group $\mathbb{F}$ on $E^{1}$, namely the subtree consisting in replacing the vertex $v$ by $1\in\mathbb{F}$, and identiying each other vertex of $K_{v}$ with the corresponding vertex in $\Gamma$.
There is a natural partial action by left translations of $\mathfrak{F_{gr}}(E)$ on the graph $\Gamma_{E}$.
Let $(E,C)$ be a separated graph. Observe that, for $v\in E^{0}$, elements in $\mathcal{X}(v)$ can be interpreted as certain finite subtrees of $K_{v}$ containing the root $v$ of the connected component $K_{v}$. Concretely, the finite subtrees $T$ which belong to $\mathcal{X}(v)$ are characterized by the property that the geodesic path between any two vertices of $T$ is a $C$-separated path (see Remark 3.8).
@@
where $T_{1}^{\prime},T_{2}^{\prime}\in\mathcal{E}$ are representatives of $T_{1}$ and $T_{2}$ respectively, such that $g_{1}\in T_{1}^{\prime}$ and $g_{2}\in T_{2}^{\prime}$. The inverse is given by $(T,g)^{-1}=((g^{-1}\cdot T^{\prime})_{0},g^{-1})$, where $T^{\prime}$ is a representative of $T$ such that $g\in T^{\prime}$.
@@
Let $(E,C)$ be a separated graph and let $\mathcal{E}$ be the semilattice of idempotents of $\mathcal{S}(E,C)$. Then there exists a (zero-preserving) isomorphism $\mathcal{Y}\cong\mathcal{E}$ given by
@@
where $I_{0}$ is the unique maximal element of $[I]$ in the semilattice order (i.e., reverse inclusion), which is characterized by the property that each maximal element of $I_{0}$ does not end in $E^{-1}$.
@@
We now analyze Examples 3.3 and 3.4 in the light of the Munn tree picture.
@@ 5. The tight spectrum
Before we continue with our analysis of (inverse) semigroups of separated graphs, let us discard a trivial part of them, namely the isolated vertices, that is, the vertices $v\in E^{0}$ with $s^{-1}(v)=r^{-1}(v)=\emptyset$. If $E^{0}_{\mathrm{iso}}$ denotes the set of isolated vertices, notice that these are idempotent elements of both semigroups $S=\tilde{\mathcal{S}}(E,C)$ or $S=\mathcal{S}(E,C)$ that are orthogonal to every other element. This means that $S$ contains a (commutative) inverse subsemigroup $E^{0}_{\mathrm{iso}}\cup\{0\}$ consisting of pairwise orthogonal idempotents that do not interact with the other non-zero elements of the semigroup. This “isolated part” of $S$ will not be interesting in any respect, as for instance in the algebras we are going to consider later, this only contributes to a trivial commutative direct sum of some copies of the underlying ground field (or ring). Hence there is no loss of generality in assuming $E^{0}_{\mathrm{iso}}=\emptyset$, so we postulate, from now on, the following:
@@
With the above notation:
@@
    The poset $\mathcal{F}$ of filters of $\mathcal{Y}$ is order-isomorphic to the poset $\mathfrak{F}$.
@@
    In particular, this order-isomorphism restricts to a bijection between the set of ultrafilters of $\mathcal{Y}$ and the subset $\mathfrak{U}\subseteq\mathfrak{F}$.
@@
We finally show that $Z$ is maximal in $\mathfrak{F}$, equivalently in $\mathfrak{F}(v)$. By way of contradiction suppose that $Z\subsetneqq\tilde{Z}$ for some $\tilde{Z}\in\mathfrak{F}(v)$. Take $g\in\tilde{Z}\setminus Z$, and let $h$ be the largest prefix of $g$ such that $h\in Z$, so that $hy\in\tilde{Z}\setminus Z$ for some $y\in E^{1}\cup E^{-1}$. Suppose first that $y\in E^{1}$. Then $y\in X$ for a unique $X\in C_{r(h)}=C_{s(y)}$. On the other hand, $h\in Z^{\prime}$ and thus $Z^{\prime}_{h}$ is a maximal local configuration. It follows that there is a unique $x\in X\cap Z^{\prime}_{h}$, so that $hx\in Z^{\prime}$. Hence $hx\in Z$, because it belongs to $Z^{\prime}$ and ends in $E^{1}$. (Observe that $h$ cannot end in $x^{-1}$, because $hy$ is a reduced word that belongs to $\tilde{Z}$, and so it belongs to $\mathbb{F}_{C}(v)$; hence $hx$ is reduced.) If $x=y$ then $hy\in Z$ which is not the case, hence $x\neq y$. But then, since $Z\subseteq\tilde{Z}$, we have $hx\in\tilde{Z}$. Since also $hy\in\tilde{Z}$, we obtain a contradiction, because $\tilde{Z}$ is $C$-compatible, and $hx$ and $hy$ are $C$-incompatible. So in any case we get a contradiction, and thus $y\in E^{-1}$. Write $y=e_{1}^{-1}$, where $e_{1}\in E^{1}$. We have $he_{1}^{-1}\in\tilde{Z}$. Moreover since $Z^{\prime}_{h}$ is a maximal local configuration, $e_{1}^{-1}\in Z^{\prime}_{h}$, which implies that $he_{1}^{-1}\in Z^{\prime}$. Now since $\tilde{Z}\in\mathfrak{F}(v)$, we can extend $he_{1}^{-1}$ to a reduced word ending in $E^{1}$, so that there are $e_{2},\dots,e_{n},d\in E^{1}$ such that
@@
    $\mathcal{N}(v):=\{x_{1}^{-1}x_{2}^{-1}\cdots x_{n}^{-1}y\in\mathbb{F}_{C}(v):n\geq 0,x_{1},\dots,x_{n},y\in E^{1}\}.$
@@
    $\mathcal{Z}(I\setminus F):=\mathcal{U}(\{I\},\{I\cup\{gg^{\prime}\}^{\downarrow}:gg^{\prime}\in F\}),$
@@
Let $(E,C)$ be a separated graph. Then the space $\mathfrak{U}$ of ultrafilters on the semilattice of idempotents of $\mathcal{S}(E,C)$ is closed in the space of all filters if and only if $(E,C)$ is finitely separated, that is, all elements of the partition $C$ are finite sets.
@@ 6.1. Leavitt path algebras and tight groupoids
of $Z$ are mutually orthogonal. Since all the sets $Z(f)$ are non-empty, $Z$ is finite, and $Y_{0}$ is infinite, we obtain a contradiction. We have thus shown that any $z\in Z$ such that $I_{z}$ contains an element $\alpha\in I_{z}\setminus I$ such that $\alpha_{I_{z}\setminus I}=\gamma_{0}e\gamma_{1}$, where $e\in Y$ and $|Y|=\infty$ is redundant. Therefore we can assume that all elements of $Z$ have the property that, for each $\alpha=\alpha_{I}\alpha_{I_{z}\setminus I}\in I_{z}$, all edges appearing (with positive exponent) in the path $\alpha_{I_{z}\setminus I}$ belong to some $X\in C^{\mathrm{fin}}$.
@@ Acknowledgments
The authors are very grateful to Ben Steinberg for his many comments and suggestions on an earlier version of the paper. In particular, the approach to Theorem 4.5 using Munn trees and the free inverse semigroup $\mathfrak{F_{is}}(E)$ is due to him. We also would like to thank the anonymous referee for their careful reading and constructive comments, which helped us improve the clarity and presentation of the paper.